Use properties of exponents to simplify:
a)100^3/2
b)-9^1/2
c)3^1/2*27^1/2
d)((1/8)^-2)^1/3
e)((121^-1)^-1/2)/22)

Answer by stanbon(75887)

Use properties of exponents to simplify:
a)100^3/2 = 10^3 = 1000
b)-9^1/2 = -3
c)3^1/2*27^(1/2) = (3*27)^(1/2) = 81^(1/2) = 9
d)((1/8)^-2)^1/3 = (1/2)^(-2) = 4
e)((121^-1)^-1/2)/22) = 11/22 = 1/2
====================
Cheers,
Stan H.